### v1

# Why Agentic Theorem Prover Works: A Statistical Provability Theory of Mathematical Reasoning Models

Abstract
Agentic theorem provers—pipelines that couple a mathematical reasoning model with library retrieval, subgoal-decomposition/search planner, and a proof assistant verifier—have recently achieved striking empirical success, yet it remains unclear which components drive performance and why such systems work at all despite classical hardness of proof search. We propose a distributional viewpoint and introduce statistical provability, defined as the finite-horizon success probability of reaching a verified proof, averaged over an instance distribution, and formalize modern theorem-proving pipelines as time-bounded MDPs. Exploiting Bellman structure, we prove existence of optimal policies under mild regularity, derive provability certificates via sub-/super-solution inequalities, and bound the performance gap of score-guided planning (greedy/top-kk/beam/rollouts) in terms of approximation error, sequential statistical complexity, representation geometry (metric entropy/doubling structure), and action-gap margin tails. Together, our theory provides a principled, component-sensitive explanation of when and why agentic theorem provers succeed on biased real-world problem distributions, while clarifying limitations in worst-case or adversarial regimes.
1 Introduction
Large language models (LLMs) have rapidly improved at mathematical reasoning (Barreto, 2026; Bryan et al., 2026; Epoch AI, 2026), and a parallel effort has produced increasingly capable agentic theorem provers: systems that combine an LLM policy, a planner for decomposition, a retriever over large libraries (such as Mathlib), and a proof assistant (such as Lean, Rocq, and Isabelle, often with SAT/SMT support) that provides rigorous feedback. Despite striking empirical progress, these pipelines are complex and highly modular, and it remains difficult to explain which components affect how—and in what sense they can overcome, or fail to overcome, the classical hardness barriers surrounding proof search. This gap between practice and theory is now a bottleneck: many design choices are guided by heuristics and ablations, while a principled understanding of provability under realistic distributions and finite compute is still missing.
Formalization itself has a long history, from the early foundations of logic (Frege) through Hilbert’s program and Gödel’s incompleteness results. It is also well known that even very simple fragments of logic already exhibit computational hardness: for example, $3$-SAT belongs to propositional logic yet is NP-complete, and decision problems for first-order logic are undecidable in general. Meanwhile, contemporary reasoning models can solve not only competition-style mathematics but also problems that look “modern” in flavor, and agentic systems can sometimes translate such reasoning into machine-checked proofs. This creates a tension. A formal proof is a discrete object—a finite sequence of symbolic transformations—so why should a stochastic generator plus a search loop succeed at all? Has the emergence of LLM-based reasoning somehow transcended complexity theory? The answer is no: classical lower bounds and undecidability results remain valid. The key question is therefore more subtle:
What changes when we replace worst-case, unbiased settings by the highly structured, biased distributions of theorem statements and proof patterns that appear in practice?
We propose the working hypothesis that resolves the apparent paradox:
Statistical provability. Real-world theorem instances are not distributed uniformly over “all statements of a given size.” Instead, they concentrate on a highly structured, repetitive subset of mathematics (libraries, idioms, canonical proof patterns). Under such biased distributions, there exist regimes where statistical methods—learned policies and learned scoring functions—can outperform generic decision procedures and brute-force search, even though unbiased/uniform formulations remain hard.
Traditional complexity theory has emphasized worst-case analysis (uniform norm over all instances) and, when considering average-case settings, often uses distributions designed to be unbiased or “flat.” In those settings, the implication “unbiased/uniform $\Rightarrow$ hard” is difficult to escape. Our focus is complementary: we ask when and why “typical/frequent $\Rightarrow$ easy” might hold, and how modern agentic pipelines exploit this. Of course, if an instance distribution were supported on worst cases, no such conclusion should be expected. The point is not to deny classical hardness, but to explain empirical success in the regimes where these systems actually operate.
1.1 Our approach: provability as a finite-horizon reachability problem
To make such statements precise, we model an agentic theorem prover as a finite-horizon Markov decision process (MDP). The state encodes the current proof obligation and verifier feedback (e.g., goals, subgoals, error messages), the action encodes the next proof step (tactic/lemma invocation, parameters, retrieval prompt, decoding controls), and the transition is induced by the proof assistant’s execution and checking. The evaluation target is time-bounded success: reach the solved state within a budget of $B$ interaction steps (or verifier calls). This leads to a canonical object:
  • the provability of an instance $x_{0}$ under budget $B$,
    $V_{B}^{*}(x_{0}):=\sup_{\pi}\Pr^{\pi}(\exists t\leq B:X_{t}\in G\mid X_{0}=x_{0}),$
    where $G$ is the goal/success set and $\pi$ ranges over policies.
We then define statistical provability as a distributional notion—provability averaged over an instance distribution $Q$:
$\mathbb{E}_{x_{0}\sim Q}[V_{B}^{\pi}(x_{0})]\quad\text{and}\quad\mathbb{E}_{x_{0}\sim Q}[V_{B}^{*}(x_{0})].$
This formalizes the central object that agentic systems implicitly optimize: not worst-case solvability, but success probability on a (highly biased) stream of problems under a fixed compute budget.
1.2 Why Bellman structure matters: certificates and component-wise effects
A major advantage of the MDP view is that it exposes the Bellman structure of the objective. This enables two kinds of results that are difficult to obtain from purely heuristic narratives:
  1. 1.
    Existence and optimality: under mild regularity (compactness/Feller-type continuity), the optimal finite-horizon value $V_{B}^{*}$ exists and is attained by a deterministic Markov policy.
  2. 2.
    Provability certificates: rather than only predicting success rates, we construct upper and lower bounds on $V_{B}^{*}$ using Bellman inequalities (sub-/super-solutions). This yields a certificate interval
    $L_{B}(x_{0})\leq V_{B}^{*}(x_{0})\leq U_{B}(x_{0}),$
    whose gap quantifies uncertainty about provability given an approximate planner/value surrogate.
To connect these certificates to learning and search, we treat the reasoning model as producing a score function (or a family of score functions) $h_{b}(x,a)$ that guides greedy selection, top-$k$ filtering, beam search, or rollouts. Deviations from the optimal value can then be controlled in terms of (i) approximation error of $h_{b}$ relative to the true one-step optimal value $Q_{b}^{*}$, (ii) statistical estimation error under adaptive data collection, and (iii) action-gap (“margin”) properties that determine when mis-ranking actually harms success. In a representative form, the analysis yields bounds where performance degrades roughly like
$\text{(effective proof length)}\times\text{(score error)}^{\text{(margin exponent)}},$
making explicit how design choices change success probability through a small set of interpretable parameters.
1.3 From heuristics to qualitative theory: what each pipeline component buys you
This framework lets us reason qualitatively (and, under assumptions, quantitatively) about the impact of common pipeline components:
  • Verifier feedback shrinks the reachable region, eliminates invalid branches early, and can increase action separations; in the theory this improves effective dimension, margins, and certificate tightness.
  • Retrieval over libraries (RAG) reduces approximation error by exposing relevant lemmas/proof schemas; in the theory this is a decrease in $\epsilon_{\mathrm{app}}$ and often a reduction of the effective horizon by enabling shortcuts.
  • Test-time search (beam/rollout) trades compute for probability mass on good trajectories; its benefit is controlled by margin tails and by how quickly near-optimal actions remain in top-$k$ sets.
  • Representation geometry (how proof states are embedded into a metric space) controls statistical rates through covering numbers / metric entropy; in favorable geometries, estimation error decays much faster with the number of verifier interactions.
1.4 Contributions
This paper develops a statistical provability theory for agentic theorem proving. Concretely:
  • MDP formalization and the notion of statistical provability. We formalize modern agentic theorem provers as finite-horizon reachability MDPs and define provability as optimal time-bounded success probability, together with its distributional counterpart over instance distributions.
  • Bellman-structure evaluation and certificates. We use Bellman recursions and inequalities to derive provability certificates and to bound the deviation between the optimal success probability and that achieved by score-guided planning algorithms.
  • A component-sensitive bound on success probability. We show how proof complexity, representation geometry (via metric entropy / doubling-type conditions), tool feedback, retrieval, test-time search, and sequential statistical complexity (e.g., sequential entropy / sequential Rademacher-type quantities) enter the provability bounds.
  • A bridge between logical hardness and statistical complexity. We position these results as a step toward connecting classical complexity barriers with the statistical structure of real mathematical corpora, clarifying when “typical” instances can be tractable for learned, agentic systems.
2 Problem Formulation
Here we formalize interactive formal theorem proving (with an LLM agent and an exact verifier/proof assistant) as a time-bounded reachability problem in a Markov decision process (MDP). We then develop a Bellman-structure–based evaluation of (i) provability (optimal success probability), (ii) deviation from optimality of score-based planning (greedy/top-kk/beam/rollout), and (iii) certificate-style upper/lower bounds. Finally, we discuss easy vs hard instances.
Notation and standing conventions
$(K,d_{K})$ denotes a compact metric space used as an embedding space for individual goals, $M(K)$ denote the set of finite positive Borel measures on $K$. For $W>0$:
$M_{\leq W}(K):=\{\mu\in M(K):\mu(K)\leq W\}.$
For a bounded function $f:K\to\mathbb{R}$:
$\|f\|_{\infty}:=\sup_{x\in K}|f(x)|,\quad\mathrm{Lip}(f):=\sup_{x\neq y}\frac{|f(x)-f(y)|}{d_{K}(x,y)},$
$\|f\|_{\mathrm{BL}}:=\max\{\|f\|_{\infty},\mathrm{Lip}(f)\}.$
The bounded Lipschitz metric on $M(K)$ is
$d_{\mathrm{BL}}(\mu,\nu):=\sup_{\|f\|_{\mathrm{BL}}\leq 1}\int_{K}fd(\mu-\nu).$
“Goal solved” is represented by the zero measure $0\in M(K)$.
All probabilities are defined on an underlying probability space on which the agent’s randomness (LLM sampling, exploration) and any environment randomness live.
3 From formal proof search to a time-bounded reachability MDP
3.1 Informal picture
At each step the agent observes a proof assistant state (goals + local context + error messages) and chooses an action (e.g., a tactic/lemma invocation, or an LLM query template with retrieval parameters). The proof assistant deterministically (or stochastically, if external tools are randomized) returns the next state. We are interested in the probability of reaching a solved state within a fixed time/step budget.
3.2 Measure-valued encoding of “a finite set of goals”
Let $\mathcal{G}_{0}$ denote the (discrete) set of syntactic goals (or goal+context objects). Fix an embedding map
$\phi:\mathcal{G}_{0}\to K.$
If a proof state $x$ contains goals $g_{1},\dots,g_{n}\in\mathcal{G}_{0}$ (allowing multiplicity), define the goal measure
$\mu(x):=\sum_{i=1}^{n}\delta_{\phi(g_{i})}\in M(K).$
We note: (1) The solved state (“no goals”) corresponds to $\mu(x)=0$. (2) The number of goals is the total mass: $\mu(x)(K)=n$. (3) If we enforce (or truncate to) $\mu(x)(K)\leq W$, then $\mu(x)\in M_{\leq W}(K)$.
Especially, we use the following compactness in Section˜4 to guarantee the existence of optimal policies via dynamic programming with maximizers.
Theorem 1 (Compactness).
If $(K,d_{K})$ is compact, then $(M_{\leq W}(K),d_{\mathrm{BL}})$ is a compact metric space.
A full proof is given in Appendix˜B.
This encoding is permutation-invariant and naturally supports variable goal counts.
3.3 MDP
We define a finite-horizon MDP:
State space: $\mathcal{X}:=M_{\leq W}(K)$ (or a truncated version; see Section˜8).
Action space: a compact metric space $(\mathcal{A},d_{\mathcal{A}})$ representing admissible control choices (tactic calls, parameters, retrieval settings, decoding strategy, etc.).
Transition kernel: $P(\cdot\mid\mu,a)$ is a Borel probability measure on $\mathcal{X}$.
Goal set: $G:=\{0\}\subset\mathcal{X}$ (closed in $d_{\mathrm{BL}}$).
We assume the proof assistant acts as a verifier: it produces correct successor states, so that the MDP’s transition is a faithful abstraction of actual interaction.
3.4 Provability as a time-bounded reachability objective
Definition 1 (Time-bounded success / provability).
Fix a horizon $B\in\mathbb{N}$. For a policy $\pi$ and initial state $\mu\in\mathcal{X}$, define the time-bounded success probability
$V_{B}^{\pi}(\mu):=\Pr^{\pi}_{\mu}\left(\exists t\in\{0,1,\dots,B\}\ \text{such that}\ \mu_{t}\in G\right).$
The optimal provability is
$V_{B}^{*}(\mu):=\sup_{\pi}V_{B}^{\pi}(\mu).$
Definition 2 (Finite-horizon value functions).
For $b=0,1,\dots,B$ define
$V_{b}^{*}(\mu):=\sup_{\pi}\Pr^{\pi}_{\mu}\left(\exists t\in\{0,1,\dots,b\}\ \text{s.t.}\ \mu_{t}\in G\right).$
Then $V_{0}^{*}(\mu)=\mathbf{1}[\mu\in G]$.
Definition 3 (One-step optimal action-value).
For $b\geq 1$ define
$Q_{b}^{*}(\mu,a):=\int_{\mathcal{X}}V_{b-1}^{*}(\mu^{\prime})P(d\mu^{\prime}\mid\mu,a).$
We have Bellman recursion (finite-horizon reachability). For $b\geq 0$:
$V_{b+1}^{*}(\mu)=\mathbf{1}[\mu\in G]\ \vee\ \sup_{a\in\mathcal{A}}Q_{b+1}^{*}(\mu,a).$
Lemma 1 (Bounds and monotonicity).
For all $b$ and $\mu$, $0\leq V_{b}^{*}(\mu)\leq 1$ and $V_{b}^{*}(\mu)\leq V_{b+1}^{*}(\mu)$.
Proof.
Immediate from the definition of reachability probabilities. ∎
4 Existence of optimal value functions and optimal Markov policies
We state mild continuity assumptions ensuring that the supremum over actions is attained.
Assumption 1 (Weak continuity / Feller-type condition).
For every bounded continuous $f:\mathcal{X}\to\mathbb{R}$, the map
$(\mu,a)\mapsto\int_{\mathcal{X}}f(\mu^{\prime})P(d\mu^{\prime}\mid\mu,a)$
is continuous on $\mathcal{X}\times\mathcal{A}$.
This implies that $(\mu_{n},a_{n})\to(\mu,a)$ entails $P(\cdot\mid\mu_{n},a_{n})\Rightarrow P(\cdot\mid\mu,a)$ weakly.
Theorem 2 (Existence of maximizers and optimal Markov policies).
Assume: (1) $\mathcal{X}=M_{\leq W}(K)$ is compact under $d_{\mathrm{BL}}$ (Theorem˜1), (2) $\mathcal{A}$ is compact, (3) the kernel $P$ satisfies Assumption˜1, (4) $G=\{0\}$ is closed (hence $\mathbf{1}[\mu\in G]$ is upper semicontinuous).
Then for each $b=1,\dots,B$: (a) The Bellman recursion defines a Borel-measurable $V_{b}^{*}$, (b) For each $\mu$, the supremum $\sup_{a\in\mathcal{A}}Q_{b}^{*}(\mu,a)$ is attained by some $a_{b}^{*}(\mu)\in\mathcal{A}$, (c) hence there exists an optimal deterministic Markov policy $\pi^{*}=(a_{b}^{*})_{b=1}^{B}$ achieving $V_{B}^{*}$.
The proof is based on backward induction. See Appendix˜C for the full proof.
5 Bellman inequalities and provability certificates
The following monotonicity principle is central: it yields upper/lower bound certificates without solving the full MDP.
Definition 4 (Bellman operator).
Define $T$ on bounded measurable $V:\mathcal{X}\to[0,1]$ by
$(TV)(\mu):=\mathbf{1}[\mu\in G]\ \vee\ \sup_{a\in\mathcal{A}}\int_{\mathcal{X}}V(\mu^{\prime})P(d\mu^{\prime}\mid\mu,a).$
Then $V_{b+1}^{*}=TV_{b}^{*}$ with $V_{0}^{*}=\mathbf{1}[\cdot\in G]$.
Lemma 2 (Monotonicity).
If $V\leq W$ pointwise then $TV\leq TW$ pointwise.
Proof.
Supremum and expectation preserve order. ∎
Theorem 3 (Sub-/super-solution certificates).
Let $(U_{b})_{b=0}^{B}$ and $(L_{b})_{b=0}^{B}$ be sequences of functions $\mathcal{X}\to[0,1]$.
If $U_{0}\geq\mathbf{1}[\cdot\in G]$ and
$U_{b+1}\geq TU_{b}\quad\text{for all }b=0,\dots,B-1,$
then $U_{b}\geq V_{b}^{*}$ for all $b$.
If $L_{0}\leq\mathbf{1}[\cdot\in G]$ and
$L_{b+1}\leq TL_{b}\quad\text{for all }b=0,\dots,B-1,$
then $L_{b}\leq V_{b}^{*}$ for all $b$.
Proof.
Induction using Lemma˜2. ∎
Definition 5 (provability certificate and certificate gap).
For an initial state $\mu_{0}$, any pair $(L_{B}(\mu_{0}),U_{B}(\mu_{0}))$ such that
$L_{B}(\mu_{0})\leq V_{B}^{*}(\mu_{0})\leq U_{B}(\mu_{0})$
is a provability certificate, and
$\mathrm{Gap}(\mu_{0}):=U_{B}(\mu_{0})-L_{B}(\mu_{0})$
is the certificate gap.
6 Score-based planning: deviation from optimal provability
We now connect LLM-driven planning to provability via score functions.
6.1 Score functions and induced policies
Let $(h_{b})_{b=1}^{B}$ be score functions
$h_{b}:\mathcal{X}\times\mathcal{A}\to[0,1].$
Define the greedy policy
$\pi_{h}(\mu,b)\in\arg\max_{a\in\mathcal{A}}h_{b}(\mu,a).$
Similarly, define top-$k$ candidate sets $\mathrm{Top}\text{-}k(h_{b}(\mu,\cdot))$ and beam-search variants (beam introduces a tree policy, but for analysis we track how often near-optimal actions remain in the candidate set).
6.2 Uniform score approximation and worst-case regret
Assumption 2 (Uniform approximation of $Q_{b}^{*}$ on a domain).
Let $\mathcal{D}\subseteq\mathcal{X}\times\mathcal{A}$ be a compact “relevant” domain (e.g., reachable pairs). Assume that for each $b$,
$\sup_{(\mu,a)\in\mathcal{D}}\big|h_{b}(\mu,a)-Q_{b}^{*}(\mu,a)\big|\leq\varepsilon_{b}.$
(How to obtain such bounds with high probability is addressed in Section˜7)
Lemma 3 (One-step suboptimality bound).
Under Assumption˜2, for every $\mu$ such that $(\mu,a)\in\mathcal{D}$ for all relevant $a$,
$\sup_{a}Q_{b}^{*}(\mu,a)-Q_{b}^{*}(\mu,\pi_{h}(\mu,b))\leq 2\varepsilon_{b}.$
Proof.
Let $a^{*}\in\arg\max_{a}Q_{b}^{*}(\mu,a)$ and $\hat{a}\in\arg\max_{a}h_{b}(\mu,a)$. Then $Q_{b}^{*}(\mu,a^{*})\leq h_{b}(\mu,a^{*})+\varepsilon_{b}\leq h_{b}(\mu,\hat{a})+\varepsilon_{b}\leq Q_{b}^{*}(\mu,\hat{a})+2\varepsilon_{b}$. Rearrange. ∎
Theorem 4 (Worst-case provability regret bound).
Under Assumption˜2, for any initial $\mu_{0}$ whose relevant trajectories remain inside $\mathcal{D}$,
$0\leq V_{B}^{*}(\mu_{0})-V_{B}^{\pi_{h}}(\mu_{0})\leq 2\sum_{b=1}^{B}\varepsilon_{b}.$
Proof.
At each step $b$, the greedy action loses at most $2\varepsilon_{b}$ in the optimal one-step value (Lemma˜3). Since finite-horizon reachability accumulates these losses additively under dynamic programming, summing yields the bound. ∎
6.3 Certificate construction from score bounds
Corollary 1 (Score-based certificate).
Under Assumption˜2, define for $b\geq 1$:
$U_{b}(\mu):=\mathbf{1}[\mu\in G]\ \vee\ \mathrm{clip}\Big(\sup_{a}h_{b}(\mu,a)+\varepsilon_{b}\Big),$
$L_{b}(\mu):=\mathbf{1}[\mu\in G]\ \vee\ \mathrm{clip}\Big(\sup_{a}h_{b}(\mu,a)-\varepsilon_{b}\Big),$
with $\mathrm{clip}(t)=\min\{1,\max\{0,t\}\}$. Then
$L_{b}(\mu)\leq V_{b}^{*}(\mu)\leq U_{b}(\mu),\qquad U_{b}(\mu)-L_{b}(\mu)\leq 2\varepsilon_{b}.$
Proof.
For each $\mu$, $\sup_{a}Q_{b}^{*}(\mu,a)\in[\sup_{a}h_{b}(\mu,a)-\varepsilon_{b},\ \sup_{a}h_{b}(\mu,a)+\varepsilon_{b}]$, then apply Theorem˜3. ∎
6.4 Top-kk/beam and a margin condition (fast rates)
Define the $(k+1)$-gap
$\Delta_{b}^{(k)}(\mu):=Q_{b}^{*}(\mu,a_{1}^{*}(\mu))-Q_{b}^{*}(\mu,a_{k+1}^{*}(\mu)),$
where $a_{i}^{*}(\mu)$ are actions sorted by $Q_{b}^{*}(\mu,\cdot)$.
Assumption 3 (Small-margin tail under occupancy).
Let $\nu_{b}$ be an occupancy distribution over states at depth $b$ (induced by the algorithm or a reference policy). Assume there exist $C_{\Delta},\beta>0$ such that for small $t$,
$\nu_{b}(\Delta_{b}^{(k)}(\mu)\leq t)\leq C_{\Delta}t^{\beta}.$
Theorem 5 (Expected fast-rate regret under margin).
If Assumptions˜2 and 3 hold, then for a top-$k$ policy (and similarly for a beam that preserves the top-$k$ set at each node),
$\mathbb{E}_{\mu_{0}\sim Q}\big[V_{B}^{*}(\mu_{0})-V_{B}^{\pi_{h,k}}(\mu_{0})\big]\ \leq\ C\sum_{b=1}^{B}\varepsilon_{b}^{\beta+1}$
for a constant $C$ depending on $C_{\Delta},\beta$.
Proof.
If $\Delta_{b}^{(k)}(\mu)>2\varepsilon_{b}$, uniform perturbation of size $\varepsilon_{b}$ cannot push the best action below rank $k$, so top-$k$ contains an optimal action. Errors only occur on $\Delta_{b}^{(k)}\leq 2\varepsilon_{b}$}, which has probability $\leq C_{\Delta}(2\varepsilon_{b})^{\beta}$, and the loss there is $O(\varepsilon_{b})$. Multiply and sum over $b$. ∎
We refer to Appendix˜D for more details on occupancy distribution, margin condition, fast rates, and what if margin fails.
7 High-probability control of approximation error: sufficient conditions
This section provides sufficient conditions under which Assumption˜2 holds with high probability, in settings that include: (1) offline pretraining from proof libraries (supervised on expert traces), (2) interactive pretraining using the verifier (rollouts, counterexamples, self-play), (3) adaptive data collection via beam search / rollouts.
7.1 Assumptions
Let $\mathcal{D}\subseteq\mathcal{X}\times\mathcal{A}$ be the compact domain on which we need uniform control.
Assumption 4 (Geometry of the relevant domain).
Assume: $\mathcal{D}$ has doubling dimension $d_{\mathcal{D}}$ (equivalently, covering numbers satisfy)
$\mathcal{N}(\rho;\mathcal{D})\leq\left(\frac{C_{\mathcal{D}}}{\rho}\right)^{d_{\mathcal{D}}}\quad(0<\rho\leq 1).$
Assumption 5 (Lipschitz score class).
Assume each hypothesis class $\mathcal{H}_{b}$ is uniformly Lipschitz on $\mathcal{D}$: $\mathcal{H}_{b}\subseteq\mathrm{Lip}_{L_{H}}(\mathcal{D},[0,1])$.
Definition 6 (Approximation error (expressivity term)).
Define
$\epsilon_{\mathrm{app},b}:=\inf_{h\in\mathcal{H}_{b}}\ \sup_{(\mu,a)\in\mathcal{D}}|h(\mu,a)-Q_{b}^{*}(\mu,a)|.$
This is a modeling term: it depends on the representation and architecture. Offline library pretraining and richer encoders can be viewed as reducing $\epsilon_{\mathrm{app},b}$ (or localizing the effective hypothesis class).
7.2 Statistical estimation error (high probability)
We now state a generic “adaptive uniform deviation” lemma sufficient for non-i.i.d. data produced by exploration.
Lemma 4 (Adaptive uniform deviation via covering numbers).
Let $(Z_{t})_{t=1}^{n}$ be a process adapted to a filtration $(\mathcal{F}_{t})$ taking values in $\mathcal{D}$. Let $\mathcal{F}$ be a class of functions $f:\mathcal{D}\to[-1,1]$. Then for any $\varepsilon>0$, with probability at least $1-\delta$,
$\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{t=1}^{n}\Big(f(Z_{t})-\mathbb{E}[f(Z_{t})\mid\mathcal{F}_{t-1}]\Big)\right|$
$\quad\leq 2\varepsilon+\sqrt{\frac{2\log\mathcal{N}(\varepsilon;\mathcal{F},\|\cdot\|_{\infty})+2\log(2/\delta)}{n}}.$
The proof is given in Appendix˜E.
Lemma 5 (Metric entropy of Lipschitz score classes on doubling domains).
There exists a constant $C$ depending on $C_{\mathcal{D}},d_{\mathcal{D}}$ such that for all $\varepsilon\in(0,1)$,
$\log\mathcal{N}(\varepsilon;\mathcal{H}_{b},\|\cdot\|_{\infty})\leq C\left(\frac{L_{H}}{\varepsilon}\right)^{d_{\mathcal{D}}}\log\left(\frac{1}{\varepsilon}\right).$
Proof.
Discretize $\mathcal{D}$ by an $\eta$-net and quantize function values; Lipschitzness controls interpolation error. Standard. ∎
Theorem 6 (A high-probability uniform approximation bound; sufficient condition).
Suppose that for each $(\mu,a)\in\mathcal{D}$ we can produce bounded targets $Y\in[0,1]$ such that
$\mathbb{E}[Y\mid\mu,a]=Q_{b}^{*}(\mu,a),$
and we fit $\hat{h}_{b}\in\mathcal{H}_{b}$ using empirical risk minimization for squared loss (or any method achieving the uniform deviation guarantees of Lemma˜4). Then with probability at least $1-\delta$,
$\sup_{(\mu,a)\in\mathcal{D}}|\hat{h}_{b}(\mu,a)-Q_{b}^{*}(\mu,a)|\leq\epsilon_{\mathrm{app},b}+\varepsilon_{\mathrm{stat}}(n,\delta),$
where (hiding polylog factors)
$\varepsilon_{\mathrm{stat}}(n,\delta)=O\left(L_{H}^{\frac{d_{\mathcal{D}}}{d_{\mathcal{D}}+2}}\left(\frac{\log(n/\delta)}{n}\right)^{\frac{1}{d_{\mathcal{D}}+2}}\right).$
The proof is given in Appendix˜F.
Remark 1 (What is $Y$).
Here $Y$ is a training dataset for hypothesis $\hat{h}$. With verifiers such as Lean and Rocq, a canonical construction is the following Monte Carlo rollout labeling: (1) Fix a horizon b1b-1. (2) Start from proof state $x$. (3) Apply action $a$ once (tactic invocation / lemma call / tool query). (4) Then follow some continuation procedure for at most b1b-1 further steps. (5) Output $Y=1$ if the proof is solved within the remaining steps, else $Y=0$.
8 Variable number of goals: truncation and tail control (compactness-preserving treatment)
The measure encoding naturally supports variable goal counts via $\mu(K)=|\text{goals}|$. However, compactness requires a mass bound.
8.1 Truncated state space and overflow event
Let $W>0$ and define $\mathcal{X}:=M_{\leq W}(K)$. Consider the overflow event
$\mathsf{Overflow}:=\left\{\max_{t\leq B}\mu_{t}(K)>W\right\}.$
In practice, $W$ can correspond to: (1) a hard cap induced by the prover (goal management), (2) a cap on maintained subgoals due to beam limits, (3) or a high-probability cap (quantile) justified empirically.
Proposition 1 (High-probability truncation).
Let $\pi$ be any policy. If
$\Pr^{\pi}_{\mu_{0}}(\mathsf{Overflow})\leq\delta_{W},$
then the true time-bounded success probability satisfies
$V_{B}^{\pi}(\mu_{0})\geq V_{B,\leq W}^{\pi}(\mu_{0})-\delta_{W},$
where $V_{B,\leq W}^{\pi}$ is the success probability in the truncated MDP that sends all overflow states to an absorbing failure state.
Proof.
Couple the true and truncated processes until the first overflow time. They agree on the complement of $\mathsf{Overflow}$. On $\mathsf{Overflow}$, truncated success is $\leq$ true success (since truncation forces failure). Hence the difference is at most $\Pr(\mathsf{Overflow})$. ∎
Consequence. All bounds proved on the compact space $M_{\leq W}(K)$ extend to the original system with an additive $\delta_{W}$ (for lower bounds) or $\delta_{W}$ (for upper bounds), depending on the direction.
9 Representation learning via certificate gap minimization
This section formalizes an objective for learning a continuous embedding/representation that makes provability certificates tight.
9.1 Embedding as a decision variable
Let $\Phi$ be a class of embeddings $\phi:\mathcal{G}_{0}\to K$ (or more generally, a class of metrics/feature maps inducing $K$ and $d_{K}$). Each $\phi$ induces: (1) a state representation $\mu_{\phi}(x)=\sum_{i}\delta_{\phi(g_{i})}\in M_{\leq W}(K)$, (2) a domain $\mathcal{D}_{\phi}\subseteq M_{\leq W}(K)\times\mathcal{A}$, (3) hypothesis classes $\mathcal{H}_{\phi,b}\subseteq\mathrm{Lip}_{L_{H}}(\mathcal{D}_{\phi},[0,1])$.
9.2 The “best certificate” within a restricted function family
Fix a function family $\mathcal{F}_{\phi}$ (e.g., all $L$-Lipschitz functions on $\mathcal{X}_{\phi}$) from which we construct sub-/super-solutions. Define the tightest certificate bounds within $\mathcal{F}_{\phi}$:
$U_{B,\phi}^{*}(\mu_{0})$
$:=\inf\left\{U_{B}(\mu_{0}):(U_{b})_{b=0}^{B}\subset\mathcal{F}_{\phi},\ U_{b+1}\geq TU_{b},\ U_{0}\geq\mathbf{1}_{G}\right\},$
$L_{B,\phi}^{*}(\mu_{0})$
$:=\sup\left\{L_{B}(\mu_{0}):(L_{b})_{b=0}^{B}\subset\mathcal{F}_{\phi},\ L_{b+1}\leq TL_{b},\ L_{0}\leq\mathbf{1}_{G}\right\}.$
These are well-defined as optimization problems over a convex cone of inequalities (when $\mathcal{F}_{\phi}$ is convex).
9.3 Certificate-gap minimization objective
A natural representation-learning objective is:
$\min_{\phi\in\Phi}\ \mathbb{E}_{\mu_{0}\sim Q}\Big[U_{B,\phi}^{*}(\mu_{0})-L_{B,\phi}^{*}(\mu_{0})\Big]\ +\ \lambda\mathsf{Comp}(\phi),$
where $\mathsf{Comp}(\phi)$ penalizes representations that are statistically difficult, e.g. via: (1) doubling dimension / covering numbers of $\mathcal{D}_{\phi}$, (2) Lipschitz constants required to represent $Q^{*}$, (3) distortion/bi-Lipschitz constraints to prevent degenerate embeddings.
Interpretation. This objective makes explicit the tradeoff: better representations reduce approximation error and increase margins (tight certificates), but overly complex geometries increase statistical error (loose certificates).
10 Discussion: Easy vs hard instances through the lens of provability bounds
The bounds in Sections˜6 and 7 show that the achievable success probability is governed by a small set of parameters: (1) effective horizon / proof length ($B$ or an effective $L_{\mathrm{eff}}$), (2) approximation error $\epsilon_{\mathrm{app}}$, (3) statistical error $\varepsilon_{\mathrm{stat}}(n)$ driven by geometry (doubling dimension, Lipschitz constants), (4) margin parameters $(C_{\Delta},\beta)$ for top-kk/beam stability, (5) overflow tail $\delta_{W}$.
A representative schematic lower bound is:
$\mathbb{E}_{\mu_{0}\sim Q}[V_{B}^{\pi_{h,k}}(\mu_{0})]$
$\quad\gtrsim\ \mathbb{E}_{\mu_{0}\sim Q}[V_{B}^{*}(\mu_{0})]-L_{\mathrm{eff}}\cdot(\epsilon_{\mathrm{app}}+\varepsilon_{\mathrm{stat}})^{\beta+1}$
$\phantom{\quad\gtrsim\ }-\delta_{W}-\text{(beam/top-$k$ bias terms)}.$
10.1 Easier instances (high success probability)
Instances tend to be easier when:
Short effective proofs: small $L_{\mathrm{eff}}$ (few steps to solve).
Large action margins: the correct next action is clearly better, yielding larger $\beta$ and smaller $C_{\Delta}$.
Low geometric complexity of reachable states: small $d_{\mathcal{D}}$ (local low-dimensional structure), improving $\varepsilon_{\mathrm{stat}}$.
Low approximation error: $Q^{*}$ is smooth w.r.t. the chosen representation and expressible by $\mathcal{H}$.
Informative verifier feedback: quickly prunes invalid branches and localizes the relevant domain $\mathcal{D}$.
10.2 Harder instances (low success probability)
Instances tend to be hard when:
Long required proofs: large $L_{\mathrm{eff}}$, causing error accumulation.
Small margins / many ambiguous states: $\Delta_{b}^{(k)}$ often near 0, making top-kk/beam fragile.
High geometric complexity: large $d_{\mathcal{D}}$ or near-discrete geometry (huge covering numbers), inflating $\varepsilon_{\mathrm{stat}}$.
Out-of-distribution (OOD): the relevant domain $\mathcal{D}$ is poorly covered by training data, inflating both $\epsilon_{\mathrm{app}}$ and $\varepsilon_{\mathrm{stat}}$.
Computationally hard subproblems embedded in the proof: if the proof requires solving tasks believed hard (e.g., certain cryptographic prediction tasks, or expensive computations not automated by libraries), then $L_{\mathrm{eff}}$ effectively explodes and margins flatten, driving success probability down under time limits.
11 Conclusion
We introduced a statistical provability framework for agentic theorem proving, treating formal proof generation as a finite-horizon reachability MDP and evaluating success by time-bounded probability rather than worst-case decidability. The Bellman structure yields both existence results and practical provability certificates, while our deviation bounds make explicit how pipeline components—verifier feedback, retrieval, test-time search, and representation geometry—shape success through approximation error, sequential complexity, and margin effects. Overall, the theory explains why modern agentic systems can be effective on biased, structured theorem distributions without contradicting classical hardness, and it highlights where limitations remain inevitable under uniform or adversarial instance regimes.
Acknowledgments
This work was supported by JST BOOST JPMJBY24E2, JST CREST JPMJCR25I5, and JSPS KAKENHI 24K21316,
Appendix A Literature Overview
This section reviews (i) learning-based theorem proving before modern LLMs, (ii) the transition to language-model-driven formal provers, (iii) recent “agentic” pipelines centered on decomposition and lemma synthesis, and (iv) emerging theory for test-time scaling (TTS) and verifier-guided inference. The goal is not an exhaustive survey but a map of the design patterns most relevant to our probabilistic analysis of end-to-end proof success.
A.1 From symbolic automation to learning-guided proof search (pre-LLM)
Before LLMs, a major theme in automated and interactive theorem proving was heuristics to guide symbolic search. Two recurring bottlenecks were (a) selecting the right premises from a large library and (b) selecting the next proof step (tactic/rule) so that search does not explode combinatorially.
TacticToe (HOL4) (Gauthier et al., 2017, 2021) learns from human proofs to predict tactics and then uses Monte Carlo tree search to explore tactic-level proof paths. In its reported experiments, TacticToe proves a substantial fraction of a standard-library benchmark under a tight CPU budget, illustrating how learned heuristics can dramatically outperform unguided search.
A second line of work developed environments and datasets that make learning-guided proving reproducible at scale. The HOList environment (HOL Light) (Bansal et al., 2019) and related systems (e.g., DeepHOL) emphasized “learning to prove” via imitation learning and reinforcement learning in a higher-order logic setting. In Lean specifically, LeanDojo (Yang et al., 2023) provides programmatic access to proof states, tactics, and premise annotations—lowering the barrier to training retrieval-augmented provers and to evaluating them under standardized conditions.
Finally, benchmark construction shaped the field’s trajectory by defining what “success” means operationally. miniF2F (Zheng et al., 2022) is influential here because it targets olympiad-style mathematics across multiple proof assistants and has become a common testbed for neural and LLM-based provers.
Relevance to this study. These pre-LLM systems already instantiate the MDP perspective: a proof state evolves under discrete actions, search is sequential/adaptive, and learned policies are evaluated by success probability under a compute budget. What changes in the LLM era is (i) the expressivity of the policy class, (ii) the role of natural-language reasoning and intermediate “thought” tokens, and (iii) the extent to which decomposition and retrieval become first-class citizens of the pipeline.
A.2 Early neural language modeling for formal proof
GPT-f (Metamath) (Polu and Sutskever, 2020) is one of the earliest studies connecting LLM to formal proof. They found new short proofs, which were accepted into the main Metamath library.
Two methodological shifts occur in this era:
  1. 1.
    From feature engineering to representation learning. Proof states and tactics can be encoded as sequences (or structured objects), letting models learn latent representations of “what to do next.”
  2. 2.
    From local decisions to global strategies. Generative models naturally produce multi-step outputs (proof sketches, tactic scripts), which encourages pipeline designs that interleave proposal, verification, and refinement.
These shifts set the stage for today’s agentic provers, where “one-step prediction” is only a component in a broader search-and-feedback loop.
A.3 Agentic theorem proving and the decomposition trend
Recent systems increasingly treat theorem proving as an agentic workflow: propose a plan, decompose into subgoals, retrieve lemmas/premises, attempt formal steps, and iterate using proof assistant feedback. A particularly visible trend is explicit subgoal decomposition / lemma synthesis—operationally analogous to introducing “cuts” in proof theory.
A few representative approaches:
  • Draft, Sketch, and Prove (DSP). (Jiang et al., 2023) DSP maps informal proofs (human-written or LLM-generated) into formal proof sketches that guide an automated prover by steering search toward easier subproblems; reported results show substantial gains on competition-style problem collections when guided by sketches.
  • DeepSeek-Prover and DeepSeek-Prover-V2. (Xin et al., 2024; Ren et al., 2025) DeepSeek-Prover focuses on data generation and training for Lean 4 theorem proving; DeepSeek-Prover-V2 explicitly emphasizes reinforcement learning for subgoal decomposition, describing a recursive pipeline that decomposes complex problems into subgoals and then synthesizes proofs of subgoals into longer reasoning traces.
  • Hilbert. (Varambally et al., 2025) Hilbert is presented as an agentic framework that aims to bridge informal reasoning and formal verification, explicitly positioning itself as combining complementary strengths of natural-language reasoning and proof assistant checking.
  • Prover Agent. (Baba et al., 2025) Prover Agent integrates LLMs with Lean and explicitly coordinates multiple roles (informal reasoning, formal proving, feedback usage), including auxiliary lemma generation.
  • Seed-Prover. (Chen et al., 2025) Seed-Prover frames itself as lemma-style whole-proof reasoning that iteratively refines proofs based on Lean feedback and intermediate lemmas, coupled with explicit test-time inference strategies enabling “deep and broad” reasoning.
  • Aristotle. (Achim et al., 2025) Aristotle is described as combining formal verification with informal reasoning and dedicated components such as geometry solvers, targeting IMO-level performance.
Unifying view. Despite varied engineering choices, these systems share a structural pattern:
  1. 1.
    Policy proposes a candidate step or a short script.
  2. 2.
    Planner introduces intermediate structure (subgoals, lemmas, proof sketches).
  3. 3.
    Retriever supplies context from a library (premises, analogous proofs).
  4. 4.
    Verifier (proof assistant kernel and/or solver) filters and provides feedback.
  5. 5.
    The loop repeats under a test-time compute budget.
This is precisely the structure we model as an MDP with an augmented agent, where success probability emerges from the interaction between (i) the base policy’s quality, (ii) decomposition effectiveness, (iii) retrieval accuracy, and (iv) the amplification provided by search and verification.
A.4 Test-time scaling and reasoning models: empirical and theoretical lenses
A second literature stream, partly orthogonal but increasingly intertwined with theorem proving, studies test-time scaling (TTS): improving performance by allocating more compute during inference, often via longer reasoning traces, repeated sampling, search, or verifier-guided selection.
A.4.1 Chain-of-thought prompting as a control knob
Zero-shot chain-of-thought prompting (“Let’s think step by step”, Kojima et al. (2022)) was popularized as a surprisingly strong baseline demonstrating that a single instruction can elicit multi-step reasoning behavior even without exemplars. From our standpoint, this is an early indication that the policy class at test time is not fixed: prompts and inference procedures can effectively change the hypothesis being deployed.
A.4.2 CoT as increased computational expressivity
Theoretical work has argued that producing intermediate reasoning tokens can expand what constant-depth transformer computations can realize, especially for tasks that are inherently serial (Feng et al., 2023). This resonates with theorem proving, where long-horizon dependencies and serial structure are fundamental: proofs are not merely long outputs, but adaptive processes where future steps depend on earlier derived subgoals.
A.4.3 Latent-variable and variance-reduction viewpoints
A complementary theoretical lens treats rationales / chains-of-thought as latent variables and studies training or inference by marginalizing over them. In particular, Phan et al. (2023) proposes maximizing marginal likelihood over rationales and discusses variance reduction in gradient estimates as models improve—connecting conceptually to self-consistency as “Monte Carlo marginalization” at inference time.
A.4.4 Which TTS strategy works depends on problem structure
Recent empirical studies argue that test-time compute does not translate into uniform gains: different tasks benefit from different allocations across sampling, self-correction, verifier-guided selection, or adaptive procedures. Wu et al. (2024) motivates one of our central modeling choices: the “difficulty” of a theorem should not be summarized solely by a single scalar (e.g., length), but also by structure—including decomposability and availability of useful lemmas in the library.
A.4.5 Verifiers can change scaling laws
Verifier-guided selection is often the mechanism that makes TTS effective in high-stakes settings. Several authors (Beirami et al., 2025; Setlur et al., 2025; Botta et al., 2025) formalize Best-of-$N$ (BoN) selection and characterize how performance depends on verifier quality. They argue for separations between verifier-based and verifier-free approaches under comparable compute/data budgets. In theorem proving, the proof assistant kernel plays the role of a near-perfect verifier for syntactic correctness, but “semantic” verifiers (e.g., premise relevance, lemma usefulness, intermediate progress) may be imperfect—so the verifier-quality dependence remains central.
A.5 Statistical theory with computational constraints
Classical computational learning theory already highlights the role of interaction and queries. A canonical example is Angluin’s $L^{\ast}$ algorithm (Angluin, 1987) for learning regular languages from membership and equivalence queries to a “teacher,” connecting sample efficiency to an interactive protocol rather than passive i.i.d. data. This viewpoint is highly suggestive for theorem proving: the proof assistant and solver feedback constitute an interactive information channel, and the agent’s objective is to adaptively extract the right information to complete a proof.
More recently, several directions attempt to incorporate computational constraints into information and generalization theory. For instance, Finzi et al. (2026) proposes “epiplexity”, an information-like quantity intended to capture what computationally bounded observers can extract from data, explicitly motivated by tensions between classical information theory and modern practice. Separately, several authors (Strobl et al., 2024) relate transformer expressivity to formal languages and circuit classes, offering a vocabulary to discuss which forms of reasoning are representationally plausible under given architectural and precision constraints.
Positioning. Our work sits at the intersection: we analyze theorem proving as an interactive sequential decision process, where success depends on both (i) computational/proof-theoretic structure (e.g., the availability of short proofs via cut/lemmas) and (ii) statistical/learning-theoretic quantities suited to adaptive interaction (sequential complexity, covering numbers). We do not claim a full unification, but we aim to put these components into a single framework where their contributions to proof success probability can be compared on the same scale.
Appendix B Proof of Theorem˜1: Compactness of $M_{\leq W}(K)$ under $d_{\mathrm{BL}}$
We prove Theorem˜1.
Lemma 6 (Arzelà–Ascoli compactness of the BL unit ball).
Let
$\mathcal{F}:=\{f:K\to\mathbb{R}:\ \|f\|_{\mathrm{BL}}\leq 1\}.$
Then $\mathcal{F}$ is compact in $(C(K),\|\cdot\|_{\infty})$.
Proof.
Functions in $\mathcal{F}$ are uniformly bounded ($\|f\|\leq 1$) and equicontinuous (Lipschitz constant $\leq 1$). Since $K$ is compact, Arzelà–Ascoli applies. ∎
Lemma 7 ($d_{\mathrm{BL}}$ metrizes weak convergence on $M_{\leq W}(K)$).
Let $(\mu_{n})$ and $\mu$ be in $M_{\leq W}(K)$. Then
$\mu_{n}\Rightarrow\mu\ \text{(weakly)}\quad\Longleftrightarrow\quad d_{\mathrm{BL}}(\mu_{n},\mu)\to 0.$
Proof.
we show
  • ($\Leftarrow$) If $d_{\mathrm{BL}}(\mu_{n},\mu)\to 0$, then $\int fd(\mu_{n}-\mu)\to 0$ for all bounded Lipschitz $f$. Bounded Lipschitz functions are dense in $C(K)$ on compact metric spaces, so convergence holds for all continuous $f$, i.e., $\mu_{n}\Rightarrow\mu$.
  • ($\Rightarrow$) Assume $\mu_{n}\Rightarrow\mu$. Suppose $d_{\mathrm{BL}}(\mu_{n},\mu)\not\to 0$. Then there exists $\varepsilon>0$ and a subsequence (still indexed by $n$) and $f_{n}\in\mathcal{F}$ such that
    $\left|\int f_{n}d(\mu_{n}-\mu)\right|\geq\varepsilon\quad\text{for all }n.$
    By Lemma A.1, $(f_{n})$ has a further subsequence converging uniformly to some $f\in\mathcal{F}$. Then
    $\left|\int f_{n}d(\mu_{n}-\mu)\right|$ $\leq\left|\int fd(\mu_{n}-\mu)\right|+\left|\int(f_{n}-f)d(\mu_{n}-\mu)\right|$
    $\leq\left|\int fd(\mu_{n}-\mu)\right|+\|f_{n}-f\|_{\infty}\cdot\|\mu_{n}-\mu\|_{\mathrm{TV}}.$
    Since $\mu_{n},\mu$ are positive with mass $\leq W$, $\|\mu_{n}-\mu\|_{\mathrm{TV}}\leq\mu_{n}(K)+\mu(K)\leq 2W$, and $\|f_{n}-f\|_{\infty}\to 0$. Also, $f$ is continuous so $\int fd(\mu_{n}-\mu)\to 0$ by weak convergence. Hence the right-hand side tends to 0, contradicting $\geq\varepsilon$.
Therefore $d_{\mathrm{BL}}(\mu_{n},\mu)\to 0$. ∎
Lemma 8 (Compactness).
$M_{\leq W}(K)$ is compact under $d_{\mathrm{BL}}$.
Proof.
$M_{\leq W}(K)$ is tight and weakly sequentially compact (by Prokhorov; on compact $K$, tightness is automatic). Any sequence has a weakly convergent subsequence; by Lemma A.2 that subsequence is Cauchy and convergent in $d_{\mathrm{BL}}$. Hence sequential compactness, hence compactness. ∎
This completes the proof of Theorem˜1.
Appendix C Proof of Theorem˜2
Proof.
We argue by backward induction on $b$.
  • Base: $V_{0}^{*}(\mu)=\mathbf{1}[\mu\in G]$ is upper semicontinuous because $G$ is closed.
  • Induction step: assume $V_{b-1}^{*}$ is upper semicontinuous and bounded. For any sequence $(\mu_{n},a_{n})\to(\mu,a)$, Assumption˜1 implies $P(\cdot\mid\mu_{n},a_{n})\Rightarrow P(\cdot\mid\mu,a)$. By the Portmanteau theorem (applied to the bounded upper semicontinuous function $V_{b-1}^{*}$),
    $\limsup_{n\to\infty}\int V_{b-1}^{*}(\mu^{\prime})P(d\mu^{\prime}\mid\mu_{n},a_{n})\leq\int V_{b-1}^{*}(\mu^{\prime})P(d\mu^{\prime}\mid\mu,a),$
    so $Q_{b}^{*}$ is upper semicontinuous on $\mathcal{X}\times\mathcal{A}$. Taking the supremum over compact $\mathcal{A}$ preserves upper semicontinuity and guarantees attainment (maximum theorem for upper semicontinuous functions on compact sets).
  • Finally, measurable selection provides a Borel measurable maximizer $\mu\mapsto a_{b}^{*}(\mu)$.
Thus optimal deterministic Markov policies exist. ∎
Appendix D Occupancy distribution, margin condition, fast rates, and what if margin fails
D.1 Occupancy distribution $\nu_{b}$: definition and meaning
In finite-horizon control, the occupancy distribution captures where the algorithm actually visits, and is the correct object for distributional assumptions (like margins).
For a (possibly randomized) policy $\pi$ and an initial distribution $Q$ over states, define the depth-$b$ state occupancy as the law of $X_{b-1}$:
$\nu_{b}^{\pi}(\cdot):=\Pr^{\pi}(X_{b-1}\in\cdot\mid X_{0}\sim Q).$
For reachability tasks, it is often more appropriate to condition on “not solved yet”:
$\tilde{\nu}_{b}^{\pi}(\cdot):=\Pr^{\pi}(X_{b-1}\in\cdot\mid X_{0}\sim Q,\ T^{\pi}\geq b),$
because decisions after termination are irrelevant.
For beam search, one can analogously define an occupancy over expanded nodes at depth $b$ induced by the beam’s node-selection rule; the key point is the same: $\nu_{b}$ is the distribution of states that matter at depth $b$ under the algorithm.
D.2 Margin condition: what it says
Let $Q_{b}^{*}(x,\cdot)$ be the optimal one-step value at depth $b$. For top-$k$ analysis, define the $(k+1)$-gap
$\Delta_{b}^{(k)}(x):=Q_{b}^{*}(x,a_{1}^{*}(x))-Q_{b}^{*}(x,a_{k+1}^{*}(x)),$
where actions are ranked by $Q_{b}^{*}$.
The margin condition (Mammen-Tsybakov-style, see e.g. Arlot and Bartlett, 2011) states that near-ties are rare:
$\nu_{b}(\Delta_{b}^{(k)}(X)\leq t)\ \leq\ C_{\Delta}t^{\beta}\quad\text{for small }t.$
Interpretation:
  • $\beta$ large $\Rightarrow$ “clear separation”: ambiguous states are rare.
  • $\beta$ small (or 0) $\Rightarrow$ “many near-ties”: small score errors often change rankings.
This is exactly the condition under which plug-in decision rules achieve fast rates in classification, transplanted here to action selection.
D.3 Proof of Theorem˜5: fast-rate proof (for the plug-in greedy policy)
To make the argument transparent, we first present the $k=1$ (greedy) case; the top-$k$ version is the same with $\Delta_{b}^{(k)}$.
Assume a uniform score approximation at depth $b$:
$\sup_{x,a\in\mathcal{D}}|h_{b}(x,a)-Q_{b}^{*}(x,a)|\leq\varepsilon.$ (1)
Let $\hat{a}(x)\in\arg\max_{a}h_{b}(x,a)$ and $a^{*}(x)\in\arg\max_{a}Q_{b}^{*}(x,a)$. Define the gap $\Delta_{b}^{(1)}(x):=Q_{b}^{*}(x,a^{*}(x))-\max_{a\neq a^{*}(x)}Q_{b}^{*}(x,a)$.
Lemma 9 (Mis-selection implies small gap).
Under Equation˜1, if $\hat{a}(x)\neq a^{*}(x)$ then $\Delta_{b}^{(1)}(x)\leq 2\varepsilon$.
Proof.
If $\hat{a}\neq a^{*}$, then
$Q_{b}^{*}(x,a^{*})\leq h_{b}(x,a^{*})+\varepsilon\leq h_{b}(x,\hat{a})+\varepsilon\leq Q_{b}^{*}(x,\hat{a})+2\varepsilon.$
Thus $Q_{b}^{*}(x,a^{*})-Q_{b}^{*}(x,\hat{a})\leq 2\varepsilon$. Since $\hat{a}$ is not optimal, the left side is at least the gap $\Delta_{b}^{(1)}(x)$, giving $\Delta_{b}^{(1)}(x)\leq 2\varepsilon$. ∎
Lemma 10 (Pointwise regret bound).
Define the one-step regret at $x$:
$r_{b}(x):=\max_{a}Q_{b}^{*}(x,a)-Q_{b}^{*}(x,\hat{a}(x)).$
Then under Equation˜1,
$r_{b}(x)\leq 2\varepsilon\ \mathbf{1}{\Delta_{b}^{(1)}(x)\leq 2\varepsilon}.$
Proof.
If $\hat{a}=a^{*}$, then $r_{b}(x)=0$. Otherwise, by Lemma˜9, $\Delta_{b}^{(1)}(x)\leq 2\varepsilon$ and also $r_{b}(x)\leq\Delta_{b}^{(1)}(x)\leq 2\varepsilon$. Combine. ∎
Now take expectation under occupancy $\nu_{b}$:
$\mathbb{E}_{\nu_{b}}[r_{b}(X)]\leq 2\varepsilon\nu_{b}(\Delta_{b}^{(1)}\leq 2\varepsilon)\leq 2\varepsilon\cdot C_{\Delta}(2\varepsilon)^{\beta}=O(\varepsilon^{\beta+1}).$
This is the fast rate: the expected regret is smaller than linear in $\varepsilon$ when $\beta>0$.
Finally, summing over depths (and propagating through Bellman recursion) yields an overall provability gap bound of the form
$\mathbb{E}[V_{B}^{*}(X_{0})-V_{B}^{\pi_{h}}(X_{0})]\ \lesssim\sum_{b=1}^{B}\mathbb{E}_{\nu_{b}}[r_{b}(X)]\ =\ O\left(B\varepsilon^{\beta+1}\right),$
up to constants and technical details about termination (see Section 4 below).
Remark 2 (Top-$k$ / beam version).
Replace $\Delta_{b}^{(1)}$ by $\Delta_{b}^{(k)}$. The same logic shows:
  • if the optimal action is pruned from top-$k$, then $\Delta_{b}^{(k)}\leq 2\varepsilon$;
  • the loss incurred by missing the optimal action is at most $\Delta_{b}^{(k)}\leq 2\varepsilon$ on that event;
  • thus expected loss scales as $\varepsilon\cdot\nu_{b}(\Delta_{b}^{(k)}\leq 2\varepsilon)=O(\varepsilon^{\beta+1})$.
D.4 If the margin condition fails: what can we still say?
If the margin tail does not decay (effectively $\beta=0$), then the above argument collapses to the worst-case linear behavior:
$\mathbb{E}_{\nu_{b}}[r_{b}(X)]\ \lesssim\ O(\varepsilon).$
More pessimistically, if there is non-negligible mass of exact ties (e.g., $\nu_{b}(\Delta=0)>0$), then ranking-based selection can be intrinsically unstable: arbitrarily small perturbations can flip the argmax on that set. In that case:
  • fast rates are not information-theoretically available from plug-in ranking alone,
  • and one must rely on search (keeping multiple candidates), additional structure (tie-breaking information, richer feedback), or accept slower rates.
Beam/top-$k$ can still help when margins fail, but then the analysis must track set coverage rather than “gap-based stability.” A common alternative is to bound success in terms of the probability that at least one of the retained actions is “good enough,” which leads to different regularity assumptions (e.g., mass of $\epsilon$-optimal actions) rather than gap tails.
Appendix E Proof of Lemma˜4 (Adaptive uniform deviation via covering numbers)
Statement. Let $(\mathcal{F}_{t})_{t\geq 0}$ be a filtration and let $Z_{t}$ be $\mathcal{F}_{t}$-measurable random variables taking values in a set $\mathcal{Z}$. Let $\mathcal{F}$ be a class of functions $f:\mathcal{Z}\to[0,1]$. For $f\in\mathcal{F}$, define the martingale difference
$D_{t}(f):=f(Z_{t})-\mathbb{E}[f(Z_{t})\mid\mathcal{F}_{t-1}].$
Fix $\varepsilon>0$ and let $\mathcal{F}_{\varepsilon}=\{f_{1},\dots,f_{M}\}$ be an $\varepsilon$-net of $\mathcal{F}$ in $\|\cdot\|_{\infty}$ (so $M=\mathcal{N}(\varepsilon;\mathcal{F},\|\cdot\|_{\infty})$). Then for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{t=1}^{n}D_{t}(f)\right|\leq 2\varepsilon+\sqrt{\frac{2\log(2M/\delta)}{n}}.$
Proof.
we show
  1. 1.
    Reduction to a finite net. For any $f\in\mathcal{F}$, choose $f_{i}\in\mathcal{F}_{\varepsilon}$ such that $\|f-f_{i}\|_{\infty}\leq\varepsilon$. Then for each $t$,
    $|D_{t}(f)-D_{t}(f_{i})|$ $=\left|\big(f(Z_{t})-f_{i}(Z_{t})\big)-\mathbb{E}\big[f(Z_{t})-f_{i}(Z_{t})\mid\mathcal{F}_{t-1}\big]\right|$
    $\leq|f(Z_{t})-f_{i}(Z_{t})|+\mathbb{E}\big[|f(Z_{t})-f_{i}(Z_{t})|\mid\mathcal{F}_{t-1}\big]$
    $\leq\varepsilon+\varepsilon=2\varepsilon.$
    Averaging over $t$ gives
    $\left|\frac{1}{n}\sum_{t=1}^{n}D_{t}(f)\right|\leq\left|\frac{1}{n}\sum_{t=1}^{n}D_{t}(f_{i})\right|+2\varepsilon.$
    Hence
    $\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{t=1}^{n}D_{t}(f)\right|\leq\max_{1\leq i\leq M}\left|\frac{1}{n}\sum_{t=1}^{n}D_{t}(f_{i})\right|+2\varepsilon.$
  2. 2.
    Concentration for each net element. For fixed $i$, the sequence $D_{t}(f_{i})$ is a martingale difference adapted to $\mathcal{F}_{t}$ and satisfies $D_{t}(f_{i})\in[-1,1]$ because $f_{i}(Z_{t})\in[0,1]$ and its conditional expectation lies in $[0,1]$. By Azuma–Hoeffding,
    $\Pr\left(\left|\sum_{t=1}^{n}D_{t}(f_{i})\right|\geq n\gamma\right)\leq 2\exp\left(-\frac{n\gamma^{2}}{2}\right).$
  3. 3.
    Union bound. By union bound over $i=1,\dots,M$,
    $\Pr\left(\max_{i\leq M}\left|\frac{1}{n}\sum_{t=1}^{n}D_{t}(f_{i})\right|\geq\gamma\right)\leq 2M\exp\left(-\frac{n\gamma^{2}}{2}\right).$
    Set the right-hand side to $\delta$ and solve:
    $\gamma=\sqrt{\frac{2\log(2M/\delta)}{n}}.$
  4. 4.
    Combine steps. With this choice of $\gamma$, with probability at least $1-\delta$,
    $\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{t=1}^{n}D_{t}(f)\right|\leq 2\varepsilon+\gamma,$
    which is the desired bound.
Appendix F Proof of Theorem˜6 (High-probability uniform approximation)
Technically, this theorem implicitly mixes two issues:
  1. 1.
    controlling estimation error under adaptive sampling, and
  2. 2.
    converting that to a uniform-in-domain sup norm bound.
F.1 Setting
  • Let $\mathcal{D}\subseteq\mathcal{X}\times\mathcal{A}$ be compact.
  • Let $Q:\mathcal{D}\to[0,1]$ denote the target function (in the main text, $Q=Q_{b}^{*}$).
  • Assume $Q$ is $L_{Q}$-Lipschitz on $\mathcal{D}$.
  • Let $\mathcal{H}\subseteq\mathrm{Lip}_{L_{H}}(\mathcal{D},[0,1])$ be the hypothesis class.
  • Define approximation error
    $\epsilon_{\mathrm{app}}:=\inf_{h\in\mathcal{H}}\ \|h-Q\|_{\infty}.$
F.2 Sampling assumption (coverage / generative access)
There exist an $\eta$-net $\{z_{1},\dots,z_{N}\}$ of $\mathcal{D}$ (with $N=\mathcal{N}(\eta;\mathcal{D})$) and, for each net point $z_{i}$, we can collect $m$ conditionally unbiased samples
$Y_{i,1},\dots,Y_{i,m}\in[0,1],\qquad\mathbb{E}[Y_{i,j}\mid z_{i}]=Q(z_{i}),$
independent across $j$ given $z_{i}$. (In practice, $Y_{i,j}$ can be a rollout success indicator from $(\mu,a)=z_{i}$.)
Let $\hat{q}_{i}:=\frac{1}{m}\sum_{j=1}^{m}Y_{i,j}$.
F.3 Constructing an estimator
Define $\hat{h}$ as a Lipschitz extension of the values $\{\hat{q}_{i}\}$ onto $\mathcal{D}$ with Lipschitz constant at most $L_{H}$ (e.g., a McShane extension). Then $\hat{h}\in\mathrm{Lip}_{L_{H}}(\mathcal{D},[0,1])$. If we further project $\hat{h}$ onto $\mathcal{H}$ (or assume $\mathcal{H}$ contains such extensions), we can ensure $\hat{h}\in\mathcal{H}$ up to $\epsilon_{\mathrm{app}}$.
F.4 Statement (Uniform sup norm bound; explicit form)
Theorem 7 (Theorem˜6, refined).
Under the setting above, with probability at least $1-\delta$,
$\|\hat{h}-Q\|_{\infty}\leq\epsilon_{\mathrm{app}}+(L_{H}+L_{Q})\eta+\sqrt{\frac{\log(2N/\delta)}{2m}}.$ (2)
In particular, if $\mathcal{D}$ has doubling dimension $d$ so that $N\lesssim(C/\eta)^{d}$ and $n:=mN$ is the total number of samples, optimizing $\eta$ yields (up to polylog factors)
$\|\hat{h}-Q\|_{\infty}\ \lesssim\epsilon_{\mathrm{app}}+(L_{H}+L_{Q})^{\frac{d}{d+2}}\left(\frac{\log(n/\delta)}{n}\right)^{\frac{1}{d+2}}.$
Proof.
Fix an $\eta$-net $\{z_{i}\}_{i=1}^{N}$ of $\mathcal{D}$. For each $i$, by Hoeffding’s inequality,
$\Pr\left(|\hat{q}_{i}-Q(z_{i})|\geq\tau\right)\leq 2\exp(-2m\tau^{2}).$
Union bound over $i=1,\dots,N$ gives: with probability at least $1-\delta$,
$\max_{i\leq N}|\hat{q}_{i}-Q(z_{i})|\leq\sqrt{\frac{\log(2N/\delta)}{2m}}.$ (3)
Now fix any $z\in\mathcal{D}$. Let $z_{i}$ be a net point with $d(z,z_{i})\leq\eta$. Then
$|\hat{h}(z)-Q(z)|$ $\leq|\hat{h}(z)-\hat{h}(z_{i})|+|\hat{h}(z_{i})-Q(z_{i})|+|Q(z_{i})-Q(z)|$
$\leq L_{H}d(z,z_{i})+|\hat{q}_{i}-Q(z_{i})|+L_{Q}d(z,z_{i})$
$\leq(L_{H}+L_{Q})\eta+\max_{j\leq N}|\hat{q}_{j}-Q(z_{j})|.$
Taking supremum over $z\in\mathcal{D}$ and using Equation˜3 yields the bound for the extension estimator. If $\hat{h}$ is then projected into $\mathcal{H}$, an additional $\epsilon_{\mathrm{app}}$ term appears by triangle inequality, giving Equation˜2. ∎
Remark 3 (Relation to Lemma˜4 / sequential complexity.).
Theorem˜7 is a clean sufficient condition relying on explicit coverage. Lemma˜4 is more general: it controls deviations under adaptive sampling and can be used to bound on-policy or occupation-weighted errors when explicit net coverage is unavailable. In this paper, both perspectives are useful:
  • net-coverage yields uniform sup norm bounds (strongest),
  • sequential Rademacher complexity (Rakhlin et al., 2015a, b; Block et al., 2021) yields distributional bounds under realistic adaptive exploration.